Please help me understand what I am doing wrong here! thanks!:
7x+3(2x+5)=10x+17
7x+6x+15=10x+17
13x+15=10x+17
-15 -15
13x=10x+2
-10 -10
ok now I am stuck! what do I do next? thank you !

Found 3 solutions by checkley71, rmromero, algebra2student:
Answer by checkley71(8403)

7X+3(2X+5)=10X+17
7X+6X+15=10X+17
13X-10X=17-15
3X=2
X=2/3
PROOF
7*2/3+3(2*2/3+5)=10*2/3+17
14/3+3(4/3+5)=20/3+17
14/3+3(4/3+15/3)=20/3+51/3
14/3+3(4+15)/3=71/3
14/3+(4+15)=71/3
14/3+19=71/3
14/3+57/3=71/3
71/3=71/3

Answer by rmromero(383)

7x+3(2x+5)=10x+17
7x+6x+15=10x+17
13x+15=10x+17
-15 -15
13x=10x+2
-10 -10 <<<<<< instead of add -10, you will add -10x



13x = 10x + 2
-10x -10x
3x = 2 <<<<<<<<< divide both side by 3 to solve for x.
x = 2/3

Answer by algebra2student(2)

What you're doing wrong is that you have to add and subtract like terms, subtracting 10 would not be a possible way.
7x+3(2x+5)=10x+17
7x+6x+15=10x+17
13x+15=10x+17
-10x -10x
3x+15=17
-15-15
3x=2
/3/3
x=2/3